OpenStudy (lgbasallote):

LGBADERIVATIVE: \[\huge y = 10^{1-x^2}\]

OpenStudy (lgbasallote):

\[\ln y = (1-x^2)\ln 10\]

OpenStudy (lgbasallote):

\[\frac 1y y' = -2x \ln 10\]

OpenStudy (lgbasallote):

\[y' = 10^{1-x^2} (-2x\ln 10)\]

OpenStudy (lgbasallote):

does that sound right?

OpenStudy (zarkon):

yes

OpenStudy (lgbasallote):

nice!

OpenStudy (helder_edwin):

perfect!